# Bayesian variable selection in sample selection models using spike-and-slab priors

### v2

@@ 5.1 Simulation scenarios
The true model is of the form in (1) with $\bm{\alpha}=\left(0.5,1,1.5,0,\ldots,0\right)/\sqrt{2}\in\mathbb{R}^{q}$, $\bm{\beta}=\left(0.25,0.5,1,0,\ldots,0\right)\in\mathbb{R}^{p}$, that is $3$ active variables and $p-3$ spurious variables, and $p=q$. The effect sizes are chosen to represent a small, medium large effect respectively. The effect sizes for the probit equation are in line with Certo et al. (2016). The intercept $\alpha_{0}$ is chosen dependent on the simulated covariates, so that the expected proportion of missing data is $0.3$. We choose $\beta_{0}=0.5$ and $\sigma=1$. We generated $n$ covariates $\bm{w}_{i}\in\mathbb{R}^{q}$ such that their marginal distributions are standard normal and $\text{cov}(w_{j},w_{k})=0.5^{\lvert j-k\rvert}$. We let $\bm{w}_{i}=\bm{x}_{i}$ so that there is no exclusion restriction. For each $\bm{w}_{i}$, we generated values of $y_{i}$ and $s_{i}$, and repeated this for $1000$ Monte Carlo replicates (keeping the covariates fixed for each scenario). We consider sample sizes $n=500$ and $1000$, and dimensions $p=10,25$ and $50$, with $\rho=0.5$. Further simulations, considering $n=2000$ and $\rho=0,0.3$ and $0.7$, can be found in Appendices F and H. We compare the following variable selection methods:
@@ 5.2 Performance measures
The evaluation metrics we use are: Sensitivity (TPR) - the proportion of active variables correctly identified; Specificity (TNR) - the proportion of inactive variables correctly identified; True model rate (TMR) - the proportion of replicates where the selected model is the true model; Model size - the average size of the selected model. For spike-and-slab samplers, the “selected model” is the model such that a variable is included if and only if the posterior inclusion probability (PIP) is greater than $0.5$. Formally, a variable $\beta_{j}$ is “included” if more than half of the posterior samples of $\gamma^{O}_{j}$ are equal to $1$, and similarly for $\alpha_{k}$.
@@ 5.3 Results
Table 1 shows the performance of each method for $n=500$ and $p=10,25$ and $50$ respectively. It should be noted that the medium and large effects are included with probability close to one, so that the sensitivity is almost entirely determined by the inclusion of the small effect. The normal spike-and-slab Gibbs sampler has the highest true model rate for all choices of $p$ considered, though the Laplace spike-and-slab is close enough in this measure of performance that it cannot be said that the normal spike-and-slab estimates the true model more often overall. The Laplace spike-and-slab includes the small effect more frequently, at the cost of a higher rate of false positives. We emphasize that these results for spike-and-slab priors depend on prior elicitation. For example, a larger slab variance for either prior distribution would improve specificity but exclude the small effect more often.
@@
Method Selection equation Outcome equation
TMR Size Sens. Spec. TMR Size Sens. Spec.
$p=10$
Normal 0.783 3.176 0.986 0.969 0.808 3.066 0.976 0.980
Laplace 0.645 3.426 0.990 0.935 0.726 3.240 0.983 0.959
$*{}^{*}$ 0.355 4.002 0.999 0.856 0.363 3.986 0.996 0.857
$*{}^{*}$ 0.558 3.458 0.982 0.927 0.762 2.919 0.945 0.988
$p=25$
Normal 0.764 3.043 0.964 0.993 0.794 2.964 0.957 0.996
Laplace 0.735 3.128 0.970 0.990 0.775 3.035 0.964 0.994
$*{}^{*}$ 0.046 6.269 0.997 0.851 0.036 6.159 0.999 0.856
$*{}^{*}$ 0.420 3.900 0.980 0.956 0.702 2.874 0.927 0.996
$p=50$
Normal 0.741 2.923 0.941 0.998 0.703 2.833 0.920 0.998
Laplace 0.720 3.003 0.947 0.997 0.698 2.914 0.931 0.997
$*{}^{*}$ 0.001 10.295 0.998 0.845 0.002 10.194 0.997 0.847
$*{}^{*}$ 0.271 4.570 0.973 0.965 0.537 2.717 0.871 0.998
$*{}^{*}$ Performance measures for Adaptive LASSO and stepwise selection were only evaluated over replicates that converged and had finite variance.
Method Selection equation Outcome equation
TMR Size Sens. Spec. TMR Size Sens. Spec.
$p=10$
Normal 0.876 3.138 1.000 0.980 0.929 3.079 1.000 0.989
Laplace 0.799 3.243 1.000 0.965 0.863 3.157 1.000 0.978
$*{}^{*}$ 0.350 4.013 1.000 0.855 0.360 4.004 1.000 0.857
$*{}^{*}$ 0.877 3.123 0.999 0.982 0.974 2.980 0.992 1.000
$p=25$
Normal 0.908 3.094 0.999 0.996 0.928 3.066 0.999 0.997
Laplace 0.871 3.145 0.999 0.993 0.896 3.105 0.999 0.995
$*{}^{*}$ 0.048 6.077 1.000 0.860 0.044 6.018 1.000 0.863
$*{}^{*}$ 0.650 3.493 1.000 0.978 0.938 3.033 0.995 0.998
$p=50$
Normal 0.919 3.077 0.999 0.998 0.943 3.041 0.997 0.999
Laplace 0.876 3.127 0.999 0.997 0.918 3.079 0.998 0.998
$*{}^{*}$ 0.002 9.855 1.000 0.854 0.001 9.625 1.000 0.859
$*{}^{*}$ 0.476 3.833 0.999 0.982 0.938 3.009 0.991 0.999
$*{}^{*}$ Performance measures for Adaptive LASSO and stepwise selection were only evaluated over replicates that converged and had finite variance.
@@ 6 Real data applications
In this section, we present two real data applications that illustrate the use of the proposed Bayesian variable selection methodology and how it compares against adaptive LASSO and stepwise selection. The first example analyzes ambulatory expenditures (Cameron and Trivedi, 2010), whereas the second example analyzes data from the RAND Health Insurance Experiment (RAND HIE), a comprehensive study on the effect of health insurance on medical expenditures. For these applications, we standardize all the covariates and examine the parameter estimates and standard deviations using the standardized data, as the true models are unknown. For the spike-and-slab methods, we additionally look at the posterior inclusion probabilities and use the same priors as in the simulation studies (Section 5), apart from using $N(0,100)$ priors on $\alpha_{0}$ and $\beta_{0}$ (to allow for large intercepts), a $Beta(1,p+q)$ prior on $r$ (to account for low sparsity) and letting $\tau_{1,\alpha}=\sqrt{3}/\pi$ and $\tau_{1,\beta}=\left(4\log(500)\right)^{-1/2}(\log(n))^{1/2}$. The samplers are run for $50,000$ iterations, with the first $5,000$ being burn-in. We report the posterior mean of $\rho$, based on the proposed spike-and-slab priors, as a Bayes’ estimate. We additionally report the results from fitting the full model using the sampleSelection R package. We present the results for Class II priors and additional results in the Appendix.
@@ 6.2 RAND data
We now analyze a data set from the RAND Health Insurance Experiment (RAND HIE). This is a comprehensive study from $1971$ to $1986$ of the impact of randomized health insurance on health care cost, utilization and outcomes. This data set was used by Cameron and Trivedi (2005) to analyze how the patient’s use of health services is affected by types of randomly assigned health insurance. In line with previous applications, we use the variable lnmeddol, the logarithm of medical expenses per individual, as the outcome variable. We let $\bm{x}$ consist of the logarithm of coinsurance rate plus 1 ($\texttt{logc}=\log(\texttt{coins}+1)$), the dummy variable for individual deductible plan (idp), the logarithm of participation incentive payment (lpi), an artificial variable fmde that is 00 if $\texttt{idp}=1$ and $\log(\max(1,\texttt{mde}/(0.01*\texttt{coins}))$ otherwise (where mde is the maximum expenditure offer), physical limitations (physlm), the number of chronic diseases (disea), dummy variables for good (hlthg), fair (hlthf) and poor (hlthp) self-rated health (where the baseline is excellent self-rated health), the log of family income (linc), the log of family size (lfam), education of household head in years (educdec), age of individual in years (xage), a dummy variable for female individuals (female), a dummy variable for individuals younger than $18$ years (child), a dummy variable for female individuals younger than $18$ years (fchild), and a dummy variable for black household heads (black). The selection variable is binexp which indicates whether the medical expenses are positive, and let $\bm{w}=\bm{x}$ so that there is no exclusion restriction. A subsample is selected so that the study year is $2$ and educdec is not “NA”. There are $5,574$ observations, of which $1,293$ have zero medical expenses so that lnmeddol is missing. Table 6 in the Appendix contains the results of each method applied to the subsample. All the methods estimate similar values for $\rho$ of around $0.73$, aside from stepwise selection which estimates $\rho=-0.03$, failing to detect any significant sample selection bias. Stepwise selection also excludes logc from the outcome, which both Adaptive LASSO and the spike-and-slab sampler include as a significant variable. The inference is otherwise similar between the normal spike-and-slab prior and Adaptive LASSO. In the selection equation, the same variables are excluded. The only difference in this regard is that the normal spike-and-slab prior excludes hlthg from the outcome, with the inclusion probability being $0.217$, whereas the Adaptive LASSO does not exclude this variable, though it does heavily shrink it to almost the same extent as the normal spike-and-slab prior. We also observe that the normal spike-and-slab prior produces larger estimates of the coefficients than Adaptive LASSO does.
@@ 6.3 Sensitivity analysis
In this section, we perform a sensitivity analysis for the spike and slab variances on the two datasets. Let $\tau_{1,\alpha}^{*}=\sqrt{3}/\pi$ and $\tau_{1,\beta}^{*}=(4\log(500))^{-1/2}(\log(n))^{1/2}$. Alongside those values, which are the defaults for the previous analysis, we consider two additional specifications. We consider $\tau_{1,\alpha}=2\tau_{1,\alpha}^{*}$ and $\tau_{1,\beta}=2\tau_{1,\beta}^{*}$, alongside the choices $\tau_{1,\alpha}=0.5\tau_{1,\alpha}^{*}$ and $\tau_{1,\beta}=0.5\tau_{1,\beta}^{*}$. In each case, the variances have been multiplied or shrunk by a factor of $4$, so that the values have change significantly. Other than the choice of spike-and-slab variances, we use the same elicitation as in the previous two sections. Table 4 shows the comparison between the three different choices on the ambulatory data. Similar tables for the spike variance and analogous tables for the RAND data can be found in Appendix I. As expected, the larger slab variances have smaller inclusion probabilities for every variable. Despite the difference in prior variances, though, the difference between the models is not drastic. The only difference in the maximum a posteriori model is the exclusion of ins in the selection equation by the larger slab variances. The posterior mean of the coefficients for variables with inclusion probability not close to $1$ are shrunk more by the larger variances, but otherwise are similar. Similar behaviour occurs for the RAND data (Table 21 in the Appendix). Due to the high collinearity between the predictors, some variables actually have higher inclusion probability under the larger slab variance, when highly correlated variables are excluded in their place. Tables 20 and 22 in the Appendix show that posterior inference based on the spike-and-slab prior is less sensitive to the spike variance than the slab variance. There is little difference in inclusion probability for most variables.
Slab variance scaling ×$\times 1$$\times 0.5$$\times 2$
PIP Est. S.D. PIP Est. S.D. PIP Est. S.D.
Selection equation
(Intercept) - 1.276 0.038 - 1.269 0.037 - 1.276 0.038
educ 1.000 0.176 0.032 1.000 0.172 0.032 0.996 0.180 0.033
age 0.949 0.108 0.039 0.977 0.108 0.034 0.914 0.108 0.044
income 0.349 0.029 0.042 0.471 0.037 0.044 0.206 0.019 0.037
female 1.000 0.320 0.031 1.000 0.319 0.030 1.000 0.320 0.031
totchr 1.000 0.603 0.055 1.000 0.586 0.052 1.000 0.604 0.055
blhisp 1.000 -0.171 0.029 1.000 -0.169 0.029 1.000 -0.174 0.029
ins 0.571 0.049 0.046 0.693 0.057 0.044 0.388 0.035 0.044
Outcome equation
(Intercept) - 6.564 0.060 - 6.562 0.057 - 6.566 0.059
educ 0.116 0.008 0.018 0.196 0.011 0.021 0.062 0.005 0.014
age 1.000 0.230 0.026 1.000 0.229 0.026 1.000 0.231 0.026
female 1.000 0.158 0.032 1.000 0.158 0.030 1.000 0.157 0.031
totchr 1.000 0.399 0.031 1.000 0.398 0.030 1.000 0.398 0.031
blhisp 0.895 -0.089 0.039 0.952 -0.092 0.033 0.844 -0.085 0.042
ins 0.033 -0.002 0.008 0.079 -0.002 0.010 0.017 -0.001 0.008
$\sigma$ 1.288 0.024 - 1.287 0.023 - 1.289 0.024
$\rho$0.265 0.154 - -0.262 0.145 - -0.271 0.150
@@ 7 Conclusion
We proposed two continuous spike-and-slab prior structures for variable selection in the context of sample selection models, and provided practical guidelines for calibrating these priors. We developed Gibbs samplers, with closed-form tractable conditionals, for sample selection models coupled with the proposed prior structures. An appealing feature of the proposed prior structures and Gibbs samplers is that they allow for using (different) scale mixture of normals for the spike and slab components. We have shown that these Gibbs samplers are scalable to the dimensions of interest in practice. Numerical results from Table 7 in Appendix E show that the spike-and-slab Gibbs sampler scales better than Adaptive LASSO and stepwise selection in higher dimensions. The time complexity is clearly not linear, so it will not scale well to dimensions $p>>200$. However, applications of interest are seldom higher dimensional than this. Our simulation study shows that the proposed Bayesian variable selection methodology offers a good performance in terms of sensitivity and specificity of the selected variables. Moreover, it exhibits similar performance to Adaptive LASSO in low dimensions and large sample sizes, while identifying the true model more frequently for extremely sparse models and smaller sample sizes. A significant advantage of our proposed methodology is that, unlike Adaptive LASSO, our proposed methodology does not suffer from convergence issues for moderately small sample sizes and higher dimensions.
There exist several natural extensions of our work. It is desirable to extend theoretical results, similar to Narisetty and He (2014) (see also Tadesse and Vannucci (2021)), to sample selection models, and to use these results to inform prior elicitation. Regarding scalability, one could extend the Skinny Gibbs algorithm (Narisetty et al., 2019) to sample selection, which would let the algorithm scale to higher dimensions, or to develop ad hoc variational Bayes methods (Tadesse and Vannucci, 2021). A similar approach could be taken with EM algorithms, such as the spike-and-slab LASSO in Ročková and George (2018). We assumed bivariate normal errors in the two-equation model (1). It would be possible to extend this model to non-normal errors, such as scale mixtures of normals (or other distributions with a tractable stochastic representation), which would allow to use the Gibbs samplers developed in Section 4 with additional steps. The proposed Gibbs samplers can also be extended to the case with binary outcomes, by using the stochastic representation of probit models through latent variables.

### v3

@@ 5.1 Simulation scenarios
The true model is of the form in (1) with $\bm{\alpha}=\left(0.5,1,1.5,0,\ldots,0\right)/\sqrt{2}\in\mathbb{R}^{q}$, $\bm{\beta}=\left(0.25,0.5,1,0,\ldots,0\right)\in\mathbb{R}^{p}$, that is $3$ active variables and $p-3$ spurious variables, and $p=q$. The effect sizes are chosen to represent a small, medium large effect respectively. The effect sizes for the probit equation are in line with Certo et al. (2016). The intercept $\alpha_{0}$ is chosen dependent on the simulated covariates, so that the expected proportion of missing data is $0.3$. We choose $\beta_{0}=0.5$ and $\sigma=1$. We generated $n$ covariates $\bm{w}_{i}\in\mathbb{R}^{q}$ such that their marginal distributions are standard normal and $\text{cov}(w_{j},w_{k})=0.5^{\lvert j-k\rvert}$. We let $\bm{w}_{i}=\bm{x}_{i}$ so that there is no exclusion restriction. For each $\bm{w}_{i}$, we generated values of $y_{i}$ and $s_{i}$, and repeated this for $1000$ Monte Carlo replicates (keeping the covariates fixed for each scenario). We consider sample sizes $n=500$ and $1000$, and dimensions $p=10,25$ and $50$, We performed simulations for $\rho=0,0.3,0.5$ and $0.7$, and found that the performance of each method was not dependent on $\rho$. For this reason, we only present the results for $\rho=0.5$ in the main text, and include results for other values of $\rho$ in Appendix G, along with results for $n=2000$ in Appendix I. We compare the following variable selection methods:
@@ 5.2 Performance measures
The evaluation metrics we use are: Sensitivity (TPR) - the proportion of active variables correctly identified; Specificity (TNR) - the proportion of inactive variables correctly identified; True model rate (TMR) - the proportion of replicates where the selected model is the true model; Model size - the average size of the selected model. For spike-and-slab samplers, the “selected model” is the median model, that is the model such that a variable is included if and only if the posterior inclusion probability (PIP) is greater than $0.5$. Formally, a variable $\beta_{j}$ is “included” if more than half of the posterior samples of $\gamma^{O}_{j}$ are equal to $1$, and similarly for $\alpha_{k}$. Barbieri and Berger (2004) shows in the normal linear case that this is the optimal predictive model.
@@ 5.3 Results
The results of the simulations can be found in Appendix F. Table 7 shows the performance of each method for $n=500$ and $p=10,25$ and $50$ respectively. It should be noted that the medium and large effects are included with probability close to one, so that the sensitivity is almost entirely determined by the inclusion of the small effect. The normal spike-and-slab Gibbs sampler has the highest true model rate for all choices of $p$ considered, though the Laplace spike-and-slab is close enough in this measure of performance that it cannot be said that the normal spike-and-slab estimates the true model more often overall. The Laplace spike-and-slab includes the small effect more frequently, at the cost of a higher rate of false positives. We emphasize that these results for spike-and-slab priors depend on prior elicitation. For example, a larger slab variance for either prior distribution would improve specificity but exclude the small effect more often.
@@ 6 Real data applications
In this section, we present two real data applications that illustrate the use of the proposed Bayesian variable selection methodology and how it compares against adaptive LASSO and stepwise selection. The first example analyzes ambulatory expenditures (Cameron and Trivedi, 2010), whereas the second example analyzes data from the RAND Health Insurance Experiment (RAND HIE), a comprehensive study on the effect of health insurance on medical expenditures. For these applications, we standardize all the covariates and examine the parameter estimates and standard deviations using the standardized data, as the true models are unknown. For the spike-and-slab methods, we additionally look at the posterior inclusion probabilities and use the same priors as in the simulation studies (Section 5), apart from using $N(0,100)$ priors on $\alpha_{0}$ and $\beta_{0}$ (to allow for large intercepts), a $Beta(1,p+q)$ prior on $r$ (to account for low sparsity) and letting $\tau_{1,\alpha}=\sqrt{3}/\pi$ and $\tau_{1,\beta}=\left(4\log(500)\right)^{-1/2}(\log(n))^{1/2}$. The samplers are run for $50,000$ iterations, with the first $5,000$ being burn-in. We report the posterior mean of $\rho$, based on the proposed spike-and-slab priors, as a Bayes’ estimate. We additionally report the results from fitting the full model using the sampleSelection R package. We present the results for Class II priors and additional results in the Appendix. We briefly consider post-selection inference. To compare models without re-fitting the data, we follow the approach of Liang et al. (2008) by comparing parameter posteriors conditional on the given models, i.e. $\left(\bm{\alpha},\bm{\beta},\sigma,\rho\right)\mid\bm{\gamma}$. In our case we consider the sub-sample where the sampled model is $\bm{\gamma}$ (and only considering parameters with $\gamma_{j}=1$). We evaluate the performance of different models using the Bayesian leave-one-out estimate of the expected log pointwise predictive density, referred to as $elpd_{loo}$ (Vehtari et al., 2017). We use the loo package in R to compute this.
@@ 6.1 Ambulatory expenditures data
For post-selection inference, we compared the $10$ most sampled models, which included the median model for spike-and-slab and the model returned by Adaptive LASSO, with results reported in 2. The best model and Adaptive LASSO include more variables than spike-and-slab, but the gain in predictive performance is relatively small, suggesting these additional variables contribute only modestly to predictive power.
$elpd_{loo}$ Model size
Best model -5086.0 12
Adaptive LASSO model -5087.5 11
Spike-and-slab median model -5087.7 10
Table 2: Values of $elpd_{loo}$ for the ten most sampled models under spike-and-slab for the ambulatory data. The $elpd_{loo}$ for each model was computed using only parameter samples corresponding to iterations where the given model was sampled. The “best model” refers to the model with the lowest $elpd_{loo}$ of the considered models.
@@ 6.2 RAND data
We now analyze a data set from the RAND Health Insurance Experiment (RAND HIE). This is a comprehensive study from $1971$ to $1986$ of the impact of randomized health insurance on health care cost, utilization and outcomes. This data set was used by Cameron and Trivedi (2005) to analyze how the patient’s use of health services is affected by types of randomly assigned health insurance. In line with previous applications, we use the variable lnmeddol, the logarithm of medical expenses per individual, as the outcome variable. We let $\bm{x}$ consist of the logarithm of coinsurance rate plus 1 ($\texttt{logc}=\log(\texttt{coins}+1)$), the dummy variable for individual deductible plan (idp), the logarithm of participation incentive payment (lpi), an artificial variable fmde that is 00 if $\texttt{idp}=1$ and $\log(\max(1,\texttt{mde}/(0.01*\texttt{coins}))$ otherwise (where mde is the maximum expenditure offer), physical limitations (physlm), the number of chronic diseases (disea), dummy variables for good (hlthg), fair (hlthf) and poor (hlthp) self-rated health (where the baseline is excellent self-rated health), the log of family income (linc), the log of family size (lfam), education of household head in years (educdec), age of individual in years (xage), a dummy variable for female individuals (female), a dummy variable for individuals younger than $18$ years (child), a dummy variable for female individuals younger than $18$ years (fchild), and a dummy variable for black household heads (black). The selection variable is binexp which indicates whether the medical expenses are positive, and let $\bm{w}=\bm{x}$ so that there is no exclusion restriction. A subsample is selected so that the study year is $2$ and educdec is not “NA”. There are $5,574$ observations, of which $1,293$ have zero medical expenses so that lnmeddol is missing. The table in Appendix 6 contains the results of each method applied to the subsample. All the methods estimate similar values for $\rho$ of around $0.73$, aside from stepwise selection which estimates $\rho=-0.03$, failing to detect any significant sample selection bias. Stepwise selection also excludes logc from the outcome, which both Adaptive LASSO and the spike-and-slab sampler include as a significant variable. The inference is otherwise similar between the normal spike-and-slab prior and Adaptive LASSO. In the selection equation, the same variables are excluded. The only difference in this regard is that the normal spike-and-slab prior excludes hlthg from the outcome, with the inclusion probability being $0.217$, whereas the Adaptive LASSO does not exclude this variable, though it does heavily shrink it to almost the same extent as the normal spike-and-slab prior. We also observe that the normal spike-and-slab prior produces larger estimates of the coefficients than Adaptive LASSO does.
Figure 1: Visualisation of a 50% credible set for the model space posterior. The black bars above variables indicate groups - if a line is not broken between two variables, they are considered to be in the same group. The numbers indicate whether a variable is included or not and the color indicates the posterior inclusion probability of each listed group permutation.
We take advantage of the recent advancements in Griffin (2024) to visualise the posterior on the model space. Figure 1 shows a 50% credible set for the model space posterior. Of most interest from this visualisation is the relation of the group consisting of lfam, xage and child, in particular the latter two. xage and child have a correlation of $-0.804$, and from Figure 1, it is clear that there is significant overlap in the information each variable brings into the model, as all three combinations involving one (or both) have similar posterior inclusion probability. In this case, both variables have inclusion probabilities greater than $0.5$, but if a higher cutoff were to be used, the visualisation suggests that at least one of these variables should be included regardless. It is also worth noting that hlthg and hlthf are grouped by this visualisation, which is where the only difference between the Adaptive LASSO model and the median model of the spike-and-slab occurs.
For post-selection inference, we consider the 20 most sampled models and use the spike-and-slab samples to obtain $elpd_{loo}$ values as in table 3. The best model in this case coincides with the model chosen by Adaptive LASSO, which includes one additional variable, but the gain in predictive performance by including this variable is modest.
$elpd_{loo}$ Model size
Adaptive LASSO model (best model) -8541.6 23
Spike-and-slab median model -8548.2 22
Table 3: Values of $elpd_{loo}$ for the 20 most sampled models under spike-and-slab for the RAND data. The $elpd_{loo}$ for each model was computed using only parameter samples corresponding to iterations where the given model was sampled. The “best model” refers to the model with the lowest $elpd_{loo}$ of the considered models.
@@ 6.3 Sensitivity analysis
In this section, we perform a sensitivity analysis for the spike and slab variances on the two datasets.
Let $\tau_{1,\alpha}^{*}=\sqrt{3}/\pi$ and $\tau_{1,\beta}^{*}=(4\log(500))^{-1/2}(\log(n))^{1/2}$. Alongside those values, which are the defaults for the previous analysis, we consider two additional specifications. We consider $\tau_{1,\alpha}=2\tau_{1,\alpha}^{*}$ and $\tau_{1,\beta}=2\tau_{1,\beta}^{*}$, alongside the choices $\tau_{1,\alpha}=0.5\tau_{1,\alpha}^{*}$ and $\tau_{1,\beta}=0.5\tau_{1,\beta}^{*}$. In each case, the variances have been multiplied or shrunk by a factor of $4$, so that the values have change significantly. Other than the choice of spike-and-slab variances, we use the same elicitation as in the previous two sections. The table 21 in Appendix J shows the comparison between the three different choices on the ambulatory data. Similar tables for the spike variance and analogous tables for the RAND data can also be found in the table in Appendix J. As expected, the larger slab variances have smaller inclusion probabilities for every variable. Despite the difference in prior variances, though, the difference between the models is not drastic. The only difference in the maximum a posteriori model is the exclusion of ins in the selection equation by the larger slab variances. The posterior mean of the coefficients for variables with inclusion probability not close to $1$ are shrunk more by the larger variances, but otherwise are similar. Similar behaviour occurs for the RAND data (the table in 23). Due to the high collinearity between the predictors, some variables actually have higher inclusion probability under the larger slab variance, when highly correlated variables are excluded in their place. The tables in 22 and 24 show that posterior inference based on the spike-and-slab prior is less sensitive to the spike variance than the slab variance. There is little difference in inclusion probability for most variables.
@@ 7 Conclusion
We proposed two continuous spike-and-slab prior structures for variable selection in the context of sample selection models, and provided practical guidelines for calibrating these priors. We developed Gibbs samplers, with closed-form tractable conditionals, for sample selection models coupled with the proposed prior structures. An appealing feature of the proposed prior structures and Gibbs samplers is that they allow for using (different) scale mixture of normals for the spike and slab components. We have shown that these Gibbs samplers are scalable to the dimensions of interest in practice. The time complexity is not linear, as clearly seen in Section Appendix A, so it will not scale well to dimensions $p>>200$. However, applications of interest studied in sample selection literature are seldom higher dimensional than this.
Our simulation study shows that the proposed Bayesian variable selection methodology offers a good performance in terms of sensitivity and specificity of the selected variables. Moreover, it exhibits similar performance to Adaptive LASSO in low dimensions and large sample sizes, while identifying the true model more frequently for extremely sparse models and smaller sample sizes. A significant advantage of our proposed methodology is that, unlike Adaptive LASSO, our proposed methodology does not suffer from convergence issues for moderately small sample sizes and higher dimensions.
There exist several natural extensions of our work. It is desirable to extend theoretical results, similar to Narisetty and He (2014), to sample selection models, and to use these results to inform prior elicitation, or use a data-driven approach to elicitate the hyperparameters instead (Tadesse and Vannucci, 2021). Regarding scalability, one could extend the Skinny Gibbs algorithm (Narisetty et al., 2019) to sample selection, which would let the algorithm scale to higher dimensions, or to develop ad hoc variational Bayes methods (Tadesse and Vannucci, 2021). A similar approach could be taken with EM algorithms, such as the spike-and-slab LASSO by Ročková and George (2018). We assumed bivariate normal errors in the two-equation model (1). It would be possible to extend this model to non-normal errors, such as scale mixtures of normals (or other distributions with a tractable stochastic representation), which would allow to use the Gibbs samplers developed in Section 4 with additional steps. The proposed Gibbs samplers can also be extended to the case with binary outcomes, by using the stochastic representation of probit models through latent variables. The focus in this paper is primarily on model selection and the problem of post-selection inference requires further investigation. Because the continuous spike-and-slab returns posterior probabilities for each model, an extension of the methodology in this paper could incorporate Bayesian model averaging (Raftery et al., 1997; Steel, 2020) into the output of the sampler directly. The issue of excluded parameters having positive posterior density away from zero under continuous spike-and-slab priors would need to be carefully considered in this context.